answer the question by writing a system of two equations and solving the system. State the answer clearly.
In an election, the winner received 430 votes more than twice as many as the lower. if there was a total of 2290 votes cast, how many did each candidate receive?

Answer by mananth(16946)

Winner ------------- x votes
loser----------------y votes
x+y=2290...............(1)
x=2y+430
x-2y=430................(2)
subtract (2) from (1)
3y=1860
/3
y= 620
Loser = 620 votes
winner = 2*620+430= 1670